For 25 army personnels, line of regression of weight of kidneys (Y) on weight of 

heart (X) is Y=0.399X+6.934 and the line of regression of weight of heart on 

weight of kidney is X−1.212Y +2.461=0. Find the correlation coefficient between 

X and Y and their mean values


1
Expert's answer
2021-06-11T09:46:00-0400

"Question:-\\
For\space 25\space army\space personnels\\
\space line\space of\space regression\space of\space weight\space of\space kidneys\space (Y)\space on\space weight\space of\space heart\space (X)\space is\\
Y=(0.399)X+6.934\\
line\space of\space regression\space of\space weight\space of\space heart\space (X)\space on\space weight\space of\space kidney\space (Y)\space is\\
X=(1.212)Y\space -2.461\\
Find\space \\
(1)the\space correlation\space coefficient\space between\space X\space and\space Y\space \\
and\\
\space (2)their\space mean\space values
\\------------------------------\\




(1)Correlation\space Coefficient\\
\space Y=0.399X+6.934\space is\space regression\space equation\space of\space y\space on\space x\\
\space 
∴b_{yx}=0.399\space ........[b_{yx}\space is\space coefficient\space of\space X\space in\space regression\space equation\space of\space y\space on\space x]\\


X=1.212Y-2.461\space is\space regression\space equation\space of\space x\space on\space y\\



∴bxy=1.212........[b_{xy}\space is\space coefficient\space of\space Y\space in\space regression\space equation\space of\space x\space on\space y]\\

Correlation\space Coefficient\space \space formula\\

r=\sqrt{byx⋅bxy}\\

r=\sqrt{0.399*1.212}\\


r=\sqrt{0.4836}\\



r=0.6954\\
-----------------------------------------
\\
(2)Mean\space of\space x\space and\space y\\
solve\space both\space equation\space and\space we\space will\space get\space mean\space of\space x\space and\space y\\
y=0.399x+6.934\\

∴y-0.399x=6.934\\

and\\
\space x-1.212y+2.461=0\\

∴x-1.21y=-2.46\\

-0.399x+y=6.934→equation(1)\\

x-1.212y=-2.461→equation(2)\\

equation(1)×1⇒-0.399x+y=6.934\\

equation(2)×0.399⇒0.399x-0.4836y=-0.9819\\

Adding\space both⇒0.5164y=5.9521\\

⇒y=\frac{5.9521}{\space 0.5164}\\
⇒y=11.5258\\

Putting\space y=11.5258\space in\space equation\space (1),\space we\space have\\

-0.399x+11.5258=6.934\\

⇒-0.399x=6.934-11.5258\\

⇒-0.399x=-4.5918\\

⇒x=\frac{4.5918}{\space 0.399}\\
\space ⇒x=11.5083\\
∴\bar{x}=11.5083\space and\space \bar{y}=11.5258\\"